Indicate in standard form the equation of the line passing through the given points.E(-2, 2), F(5, 1)

Answer by Cromlix(4381)

Using y2 - y1/x2 - x1 to find the gradient
1 - 2/5 -(-2)
-1/7
Using y - b = m(x - a) with F (5,1)
y - 1 = -1/7(x - 5)
y - 1 = -1/7x + 5/7
y = -1/7x + 5/7 + 7/7
= y = -1/7x + 12/7
OR
= 7y = -x + 12
Hope this helps.
:-)